# Future and AI-Ready Data Strategies

### v1

@@ Introduction
In 2022, the Open Government Data Act was issued. The Act mandates all federal agencies to “publish their information online as open data, using standardized, machine-readable data formats, with their metadata included in the Data.gov catalog.” Then, 2023 became the Year of Open Science. Notably, the NIH’s Data Management and Sharing Policy went into effect at the start of the year, and the NSF invested \$12.5 million into their Findable, Accessible, Interoperable, and Reusable Open Science Research Coordination Networks (FAIROS-RCN) program [^3] https://crsreports.congress.gov/product/pdf/IF/IF12299.
@@
Publishing DOC datasets on public repositories would make them easily searchable and discoverable for developers. Platforms like Hugging Face, GitHub, Google, and Meta Research’s Papers with Code are engineered for searchability, navigation, and discoverability, where developers can find the datasets most useful for their use cases. The most notable platform is GitHub. GitHub has logged collaborative version control so users can share their code and data repositories and see logged changes. The community can also contribute to repositories, while the owner retains control over which contributions are accepted. [^8] https://huggingface.co/docs/hub/en/datasets-adding Hugging Face, the largest ML dataset host and built-in search engine, would help DOC datasets be found on its platform and in public search engine results. Additionally, making DOC assets available in public data repositories would help them be seen by those who need them. One is the Hugging Face data repository, which has metadata UI and dataset cards. Their metadata support lets developers add information about their dataset, including its licensing. [^9] https://paperswithcode.com/datasets. Papers with Code is a similar platform and another highly-frequented webpage for datasets.

### v2

@@ Introduction
In 2022, the Open Government Data Act was issued. The Act mandates all federal agencies to “publish their information online as open data, using standardized, machine-readable data formats, with their metadata included in the Data.gov catalog. Then, 2023 became the Year of Open Science. Notably, the NIH’s Data Management and Sharing Policy went into effect at the start of the year, and the NSF invested \$12.5 million into their Findable, Accessible, Interoperable, and Reusable Open Science Research Coordination Networks (FAIROS-RCN) program. [^3] https://crsreports.congress.gov/product/pdf/IF/IF12299
@@
Publishing DOC datasets on public repositories would make them easily searchable and discoverable for developers. Platforms like Hugging Face, GitHub, Google, and Meta Research’s Papers with Code are engineered for searchability, navigation, and discoverability, where developers can find the datasets most useful for their use cases. The most notable platform is GitHub. GitHub has logged collaborative version control so users can share their code and data repositories and see logged changes. The community can also contribute to repositories, while the owner retains control over which contributions are accepted. [^8] https://huggingface.co/docs/hub/en/datasets-adding Hugging Face, the largest ML dataset host and built-in search engine, would help DOC datasets be found on its platform and in public search engine results. Additionally, making DOC assets available in public data repositories would help them be seen by those who need them. One is the Hugging Face data repository, which has metadata UI and dataset cards. Their metadata support lets developers add information about their dataset, including its licensing. [^9] https://paperswithcode.com/datasets Papers with Code is a similar platform and another highly-frequented webpage for datasets.